# Safe Model-Based Multi-Agent Mean-Field Reinforcement Learning

### v1

@@ Appendix C Implementation Details
In this section, we provide additional details on the practical implementation of Safe-$\text{M}^{3}$-UCRL. In particular, we discuss Probabilistic Ensemble model [13, 30] to implement the statistical model from Section 3.2 in Section C.1, the hallucinated control reparametrization from Section 4 in more detail in Section C.2, and optimization methods to solve Section 4 in Section C.3
C.1 Probabilistic Ensemble Model of Transition Dynamics
As discussed in Section 3.2, we take a model-based approach to handling the unknown dynamics $f$. The representative agent learns the Statistical Model (see Section 3.2) of the dynamics from the observed transitions $\cup_{i=1}^{n-1}\mathcal{D}_{i}$ at the beginning of each episode $n$, where $\mathcal{D}_{i}=\{(z_{i,t},s_{i,t+1})\}_{t=0}^{T-1}$ with $z_{i,t}=(s_{i,t},\mu_{i,t},a_{i,t})$. We use Probabilistic Ensemble [13, 30] that consist of $K$ neural networks parametrized by $\theta_{k}$ for $k\in\{1,\ldots,K\}$ (the episode index should be clear from the context so we omit it for the notation simplicity). Each neural network $f_{\theta_{k}}$ returns a mean vector, $\bm{m}_{\theta_{k}}(z)\in\mathcal{S}\subseteq\mathbb{R}^{p}$, and a covariance matrix $\bm{\Sigma}_{\theta_{k}}(z)\in\mathbb{R}^{p\times p}$, that represents the aleatoric uncertainty. We further assume diagonal covariance matrices. These outputs then form Gaussian distributions from which new states are sampled, i.e., $s_{t+1}\sim\mathcal{N}(\bm{m}_{\theta_{k}}(z_{t}),\bm{\Sigma}_{\theta_{k}}(z_{t}))$ for $t=0,\ldots,T-1$. The models are trained with the negative log-likelihood loss function (NLL), $L(\theta)=-\sum_{\mathcal{D}_{1:n-1}}\log\mathbb{P}(s_{t+1}|z_{t})$ as described in [30]. The ensemble means and the aleatoric and epistemic uncertainties are then estimated as follows:
$\bm{m}_{n-1}(\cdot)$ $=\frac{1}{K}\sum_{k=1}^{K}\bm{m}_{\theta_{k}}(\cdot),$
$\bm{\Sigma}_{n-1}^{e}(\cdot)$ $=\frac{1}{K-1}\sum_{k=1}^{K}(\bm{m}_{\theta_{k}}(\cdot)-\bm{m}_{n-1}(\cdot))(\bm{m}_{\theta_{k}}(\cdot)-\bm{m}_{n-1}(\cdot))^{\top},$
@@ C.1 Probabilistic Ensemble Model of Transition Dynamics
where $\bm{m}_{n-1}(\cdot)$ is the ensemble prediction for the mean, while $\bm{\Sigma}_{n-1}^{a}(\cdot)$ and $\bm{\Sigma}_{n-1}^{e}(\cdot)$ denote the aleatoric and epistemic uncertainty estimates, respectively. Note that the epistemic estimate $\bm{\Sigma}_{n-1}^{e}(\cdot)$ is used to construct the calibrated model (see Section 3.2 and Assumption 4).
Even though Gaussian Processes (GPs) are proven to be calibrated under certain regularity assumptions [1, 50], we chose probabilistic ensemble model due to their better practical performance, i.e., scalability to higher dimensions and larger datasets. The disadvantage of the probabilistic ensemble is that, unlike GPs, it does not guarantee the calibrated model (see Assumption 4). Nevertheless, it can be recalibrated using one-step ahead predictions and temperature scaling as shown in [36]. We note that in our experiment (see Section 5), such recalibration was not needed, and the above-defined mean and epistemic uncertainty outputs were sufficiently accurate for training Safe-$\text{M}^{3}$-UCRL.
C.2 Hallucinated Control Implementation Trick
In Section 4, we introduced Safe-$\text{M}^{3}$-UCRL, a model that optimizes over the confidence set of dynamics functions $\mathcal{F}_{n-1}$ and admissible policy profiles $\Pi$ (see Section 4). Unfortunately, optimizing directly over the function space is usually intractable since $\mathcal{F}_{n-1}$ is not convex, in general, [16]. Thus, to make the optimization tractable, we describe a hallucinated control trick, which leads to a practical reformulation (see Section 4). The structure in $\mathcal{F}_{n-1}$ allows us to parametrize the problem and use gradient-based optimization to find a policy profile $\bm{\pi}_{n}^{*}$ at every episode $n$. Namely, we use a mean-field variant of an established approach known as Hallucinated Upper Confidence Reinforcement Learning (H-UCRL) [15, 39, 43]. We introduce an auxiliary function $\eta:\mathcal{Z}\rightarrow[-1,1]^{p}$, where $p$ is the dimensionality of the state space $\mathcal{S}$, to define hallucinated dynamics model:
@@ C.2 Hallucinated Control Implementation Trick
where $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ are estimated from the past observations collected until the end of the previous episode $n-1$. Notice that $\tilde{f}_{n-1}$ is calibrated for any $\eta(\cdot)$ under Assumption 4, i.e., $\tilde{f}_{n-1}\in\mathcal{F}_{n-1}$. Assumption 4 further guarantees that every function $\tilde{f}_{n-1}$ can be expressed in the auxiliary form in Equation 9:
@@
$\forall\tilde{f}_{n-1}\in\mathcal{F}_{n-1}\;\exists\eta:\mathcal{Z}\rightarrow[-1,1]^{p}\;\text{such that}\;\tilde{f}_{n-1}(z)=\bm{m}_{n-1}(z)+\beta_{n-1}\bm{\Sigma}_{n-1}(z)\eta(z),\;\forall z\in\mathcal{Z}.$
@@
Furthermore, note that, for a fixed individual policy $\pi$, the auxiliary function $\eta(z)=\eta(s,\mu,\pi(s,\mu))=\eta(s,\mu)$ has the same functional form as the policy $\pi$. This turns $\eta(\cdot)$ into a policy that exerts hallucinated control over the epistemic uncertainty of the confidence set of dynamics functions $\mathcal{F}_{n-1}$ [15]. The reformulation of the optimization problem in Section 4 allows us to optimize over parametrizable functions (e.g., neural networks) $\bm{\pi},\eta(\cdot),\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ using gradient-based methods on the functions’ parameters. Note that the parametrization of $\eta(\cdot)$ must be sufficiently flexible not to restrict the space of $\tilde{f}_{n-1}$. In Section C.3, we provide several algorithms that can be used to solve this optimization problem.
C.3 Optimization Methods (MF-BPTT & MF-DDPG)
@@
In this section, we describe two algorithms to solve the optimization problem in Section 4. Namely, in Section C.3.1, we outline the key steps to apply the mean-field variant of the Back Propagation Through Time (BPTT) when a differentiable simulator is available and, in Section C.3.2, we describe the mean-field variant of the Deep Deterministic Policy Gradient (DDPG) [34] algorithm.
C.3.1 Mean-Field Back-Propagation-Through-Time (MF-BPTT)
@@
Mean-Field Back-Propagation-Through-Time (MF-BPTT) assumes access to a differentiable simulator that returns a policy rollout given transition and policy functions. In each episode $n$, the representative agent initializes the policy profile $\bm{\pi}_{n}$, the auxiliary function $\eta(\cdot)$, and the estimated dynamics $\tilde{f}_{n-1}$ using the mean vector $\bm{m}_{n-1}(\cdot)$ and covariance matrix $\bm{\Sigma}_{n-1}(\cdot)$ according to Equation 7. Then, the representative agent repeatedly calls the simulator with inputs $\bm{\pi}_{n}$ and $\tilde{f}_{n-1}$ that returns the episode reward defined in Equation 8a as a differentiable object. After each policy rollout, a gradient ascent step is carried out on the parameters of $\bm{\pi}_{n}$ and $\eta(\cdot)$ before calling the simulator again. To simplify the notation, we overload the notation $\bm{\pi}_{n}^{\psi}$ with the combination of the two policy functions $\bm{\pi}_{n}$ and $\eta(\cdot)$, i.e., $\bm{\pi}_{n}^{\psi}=(\bm{\pi}_{n}^{\psi},\eta^{\psi})$ where the superscript $\psi$ represent the parameters of both functions. We outline the described steps in Algorithm 2. During the optimization phase (Line 3 in Algorithm 1), we optimize both functions jointly. However, during the execution (Line 4 in Algorithm 1), we only use the outputs corresponding to the policy profile $\bm{\pi}_{n}$. The main distinction in Algorithm 2 compared to traditional BPTT lies inside the simulator that has to simulate the mean-field distribution $\tilde{\mu}_{n,t}$ for $t=0,\dots,T$ for each parameter update and calculate gradients with respect to this time dependency as well. More details on the implementation used for our experiments reported in Section 5 are provided in Appendix D.
@@
1:Safety constraint $h_{C}(\cdot)$, calibrated dynamics $\tilde{f}_{n-1}$ represented by $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$, initial mean-field distribution $\mu_{0}$, known reward $r(\cdot)$, constants $C_{n,t}$, safety Lipschitz constant $L_{h}$; hyperparameter $\lambda$, number of epochs $K$, number of rollout steps $T$
@@
4:     Initialize $\tilde{\mu}_{0}\leftarrow\mu_{0}$, $\tilde{s}_{0}\sim\mu_{0}$
@@
7:         $\tilde{a}_{t}\leftarrow\pi^{\psi}_{t}(\tilde{s}_{t},\tilde{\mu}_{t})$
8:         $\tilde{s}_{t+1}\leftarrow\tilde{f}_{n-1}(\tilde{s}_{t},\tilde{\mu}_{t},\tilde{a}_{t})+\varepsilon_{t}$
@@
11:     Update $\psi$ with a gradient ascent:
12:     $\quad\;\,\nabla_{\psi}\sum_{t=0}^{T-1}r(\tilde{s}_{t},\tilde{\mu}_{t},\tilde{a}_{t})+\lambda\log(h_{C}(\tilde{\mu}_{t+1})-L_{h}C_{n,t+1})$
@@
C.3.2 Mean-Field Deep Deterministic Policy Gradient (MF-DDPG)
In this section, we adopt the Deep Deterministic Policy Gradient (DDPG) algorithm [34] to the mean-field control problem. DDPG is a model-free actor-critic algorithm, hence, it can optimize Section 4 without the assumption of a differentiable simulator. However, it can not be applied directly to the problem because the Q-value for $\tilde{z}_{n,t}=(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})$ is ambiguous. The important insight here is that the value of a certain state $\tilde{s}_{n,t}$ of the environment reflects the whole population, i.e., the expected reward over the remainder of an episode for a given $\tilde{\mu}_{n,t}$ if every agent in every state $\tilde{s}_{n,t}$ chooses actions following $\pi_{n,t}$. In essence, the Q-value is a function of $\tilde{\mu}_{n,t}$ and $\pi_{n,t}$ and not $\tilde{z}_{n,t}$.
To overcome this issue, we introduce the lifted mean-field Markov decision process (MF-MDP) similarly to [9, 22, 23, 41, 43]. First, we rewrite the reward function as a function of the mean-field distribution and the policy, i.e.,
$\tilde{r}(\tilde{\mu}_{n,t},\pi_{n,t})=\int_{\mathcal{S}}r(s,\tilde{\mu}_{n,t},\pi_{n,t}(s,\tilde{\mu}_{n,t}))\tilde{\mu}(ds).$
To simplify the notation, we overload the notation $\bm{\pi}_{n}^{\psi}=(\bm{\pi}_{n}^{\psi},\eta^{\psi})$ as described in Section C.3.1. Then, we restate Section 4 as:

### v2

@@ Appendix C Implementation Details
In this section, we provide additional details on the practical implementation of Safe-M3-UCRL. In particular, we discuss Probabilistic Neural Network Ensemble model Chua et al. (2018); Lakshminarayanan et al. (2017) to implement the statistical model from Section 3.3 in Section C.1, the hallucinated control reparametrization from Section 4 in more detail in Section C.2, and optimization methods to solve Section 4 in Section C.3
C.1. Probabilistic Neural Network Ensemble Model of Transitions
As discussed in Section 3.3, we take a model-based approach to handling unknown transitions $f$. The representative agent learns the Statistical Model (see Section 3.3) of the transitions from the observed transitions $\cup_{i=1}^{n-1}\mathcal{D}_{i}$ at the beginning of each episode $n$, where $\mathcal{D}_{i}=\{(z_{i,t},s_{i,t{}+{}1})\}_{t=0}^{T-1}$ with $z_{i,t}=(s_{i,t},\mu_{i,t},a_{i,t})$. We use Probabilistic Neural Network Ensemble Chua et al. (2018); Lakshminarayanan et al. (2017) that consists of $K$ neural networks parametrized by $\theta_{k}$ for $k\in\{1,\ldots,K\}$ (the episode index should be clear from the context so we omit it for the notation simplicity). Each neural network $f_{\theta_{k}}$ returns a mean vector, $\bm{m}_{\theta_{k}}(z)\in\mathcal{S}\subseteq\mathbb{R}^{p}$, and a covariance function $\bm{\Sigma}_{\theta_{k}}(z)\in\mathbb{R}^{p\times p}$, that represents the aleatoric uncertainty. We further assume diagonal covariance functions. These outputs then form Gaussian distributions from which new states are sampled, i.e., $s_{t{}+{}1}\sim\mathcal{N}(\bm{m}_{\theta_{k}}(z_{t}),\bm{\Sigma}_{\theta_{k}}(z_{t}))$ for $t{}={}0,\ldots,T-1$. The models are trained with the negative log-likelihood loss function (NLL), $L(\theta){}={}-\sum_{\mathcal{D}_{1:n-1}}\log\mathbb{P}(s_{t{}+{}1}|z_{t})$ as described in Lakshminarayanan et al. (2017). The ensemble means, and the aleatoric and epistemic uncertainties are then estimated as follows
$\bm{m}_{n-1}(\cdot)$ ${}={}\frac{1}{K}\sum_{k=1}^{K}\bm{m}_{\theta_{k}}(\cdot)$
$\bm{\Sigma}_{n-1}^{e}(\cdot)$ ${}={}\frac{1}{K-1}\sum_{k=1}^{K}(\bm{m}_{\theta_{k}}(\cdot)-\bm{m}_{n-1}(\cdot))(\bm{m}_{\theta_{k}}(\cdot)-\bm{m}_{n-1}(\cdot))^{\top}$
@@ C.1. Probabilistic Neural Network Ensemble Model of Transitions
where $\bm{m}_{n-1}(\cdot)$ is the ensemble prediction for the mean, while $\bm{\Sigma}_{n-1}^{a}(\cdot)$ and $\bm{\Sigma}_{n-1}^{e}(\cdot)$ denote the aleatoric and epistemic uncertainty estimates, respectively. Note that the epistemic uncertainty estimate $\bm{\Sigma}_{n-1}^{e}(\cdot)$ is used to construct the calibrated model (see Section 3.3 and Assumption 4).
Even though Gaussian Processes (GPs) are proven to be calibrated under certain regularity assumptions Abbasi-Yadkori and Szepesvaŕi (2011); Srinivas et al. (2010), we chose probabilistic neural network ensemble model due to their better practical performance, i.e., scalability to higher dimensions and larger datasets. The disadvantage of the probabilistic neural network ensemble is that, unlike GPs, it does not guarantee the calibrated model (see Assumption 4). Nevertheless, it can be recalibrated using one-step ahead predictions and temperature scaling as shown in Malik et al. (2019). We note that in our experiment (see Section 5), such recalibration was not needed, and the above-defined mean and epistemic uncertainty outputs were sufficiently accurate for training Safe-M3-UCRL (see Figure 10(c)).
C.2. Hallucinated Control Implementation Trick
In Section 4, we introduced Safe-M3-UCRL, a model that optimizes over the confidence set of transitions $\mathcal{F}_{n-1}$ and admissible policy profiles $\Pi$ (see Section 4). Unfortunately, optimizing directly over the function space is usually intractable since $\mathcal{F}_{n-1}$ is not convex, in general, Dani et al. (2008). Thus, to make the optimization tractable, we describe a hallucinated control trick, which leads to a practical reformulation (see Section 4). The structure in $\mathcal{F}_{n-1}$ allows us to parametrize the problem and use gradient-based optimization to find a policy profile $\bm{\pi}_{n}^{*}$ at every episode $n$. Namely, we use the mean-field variant of an established approach known as Hallucinated Upper Confidence Reinforcement Learning (H-UCRL) Curi et al. (2020); Moldovan et al. (2015); Pásztor et al. (2023). We introduce an auxiliary function $\eta:\mathcal{Z}\rightarrow[-1,1]^{p}$, where $p$ is the dimensionality of the state space $\mathcal{S}$, to define hallucinated transitions
@@ C.2. Hallucinated Control Implementation Trick
where $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ are estimated from the past observations collected until the end of the previous episode $n-1$. Notice that $\tilde{f}_{n-1}$ is calibrated for any $\eta(\cdot)$ under Assumption 4, i.e., $\tilde{f}_{n-1}\in\mathcal{F}_{n-1}$. Assumption 4 further guarantees that every function $\tilde{f}_{n-1}$ can be expressed in the auxiliary form in Equation 11
@@
$\begin{split}&\forall\tilde{f}_{n-1}\in\mathcal{F}_{n-1}\;\exists\eta:\mathcal{Z}\rightarrow[-1,1]^{p}\;\text{such that}\\ {}&\tilde{f}_{n-1}(z){}={}\bm{m}_{n-1}(z){}+{}\beta_{n-1}\bm{\Sigma}_{n-1}(z)\eta(z),\;\forall z\in\mathcal{Z}.\end{split}$
@@
Furthermore, note that, for a fixed individual policy $\pi$, the auxiliary function $\eta(z){}={}\eta(s,\mu,\pi(s,\mu)){}={}\eta(s,\mu)$ has the same functional form as the policy $\pi$. This turns $\eta(\cdot)$ into a policy that exerts hallucinated control over the epistemic uncertainty of the confidence set of transitions $\mathcal{F}_{n-1}$ Curi et al. (2020). The reformulation of the optimization problem in Section 4 allows us to optimize over parametrizable functions (e.g., neural networks) $\bm{\pi}$ and $\eta(\cdot)$ using gradient-based methods on the functions’ parameters. Notice that the shared functional form of $\bm{\pi}$ and $\eta(\cdot)$ allows us to conveniently represent them by a single neural network. Further, note that the parametrization of $\eta(\cdot)$ must be sufficiently flexible not to restrict the space of $\tilde{f}_{n-1}$. In Section C.3, we provide several algorithms that can be used to solve this optimization problem.
C.3. Optimization Methods – MF-BPTT and MF-DDPG
@@
In this section, we describe two algorithms to solve the optimization problem in Section 4. Namely, in Section C.3.1, we outline the key steps to apply the mean-field variant of the Back-Propagation-Through-Time (BPTT) when a differentiable simulator is available and, in Section C.3.2, we describe the mean-field variant of the Deep Deterministic Policy Gradient (DDPG) Lillicrap et al. (2015) algorithm appropriate for non-differentiable simulators.
C.3.1. Mean-Field Back-Propagation-Through-Time (MF-BPTT)
@@
Mean-Field Back-Propagation-Through-Time (MF-BPTT) assumes access to a differentiable simulator that returns a policy rollout given transition and policy functions. In each episode $n$, the representative agent initializes the policy profile $\bm{\pi}_{n}$, the auxiliary function $\eta(\cdot)$, and the estimated transitions $\tilde{f}_{n-1}$ using the mean $\bm{m}_{n-1}(\cdot)$ and covariance $\bm{\Sigma}_{n-1}(\cdot)$ functions according to Equation 8. Then, the representative agent repeatedly calls the simulator with inputs $\bm{\pi}_{n}$ and $\tilde{f}_{n-1}$ that returns the episode reward defined in Equation 9a as a differentiable object. After each policy rollout, a gradient ascent step is carried out on the parameters of $\bm{\pi}_{n}$ and $\eta(\cdot)$ before calling the simulator again. To simplify the notation, we overload the notation $\bm{\pi}_{n}^{\psi}$ with the combination of the two policy functions $\bm{\pi}_{n}$ and $\eta(\cdot)$, i.e., $\bm{\pi}_{n}^{\psi}{}={}(\bm{\pi}_{n}^{\psi},\eta^{\psi})$ where the superscript $\psi$ represent the parameters of both functions. We outline the described steps in Algorithm 2. During the optimization phase (Line 3 in Algorithm 1), we optimize both functions, $\bm{\pi}_{n}$ and $\eta(\cdot)$, jointly. However, during the execution (Line 4 in Algorithm 1), we only use the outputs corresponding to the policy profile $\bm{\pi}_{n}$. The main distinction in Algorithm 2 compared to traditional BPTT lies inside the simulator that has to simulate the mean-field distribution $\tilde{\mu}_{n,t}$ for $t=1,\dots,T$ for each parameter update and calculate gradients with respect to this time dependency as well. More details on the implementation used for our experiments reported in Section 5 are provided in Section D.3 and Section E.2.
@@
1:Safety constraint $h_{C}(\cdot)$, calibrated transitions $\tilde{f}_{n-1}$ represented by $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$, initial mean-field distribution $\mu_{0}$, known reward $r(\cdot)$, constants $C_{n,t}$, safety Lipschitz constant $L_{h}$; hyperparameter $\lambda$, number of epochs $K$, number of rollout steps $T$
@@
4:     Initialize $\tilde{\mu}_{0}\leftarrow\mu_{0}$, $\thinspace{}_{0}\sim\mu_{0}$
@@
7:         $\tilde{a}_{t}\leftarrow\pi^{\psi}_{t}(\thinspace_{t},\tilde{\mu}_{t})$
8:         $\thinspace{}_{t{}+{}1}\leftarrow\tilde{f}_{n-1}(\thinspace_{t},\tilde{\mu}_{t},\tilde{a}_{t}){}+{}\varepsilon_{t}$
@@
11:     Update $\psi$ with gradient ascent
12:     $\quad\;\,\nabla_{\psi}\sum_{t=0}^{T-1}r(\thinspace_{t},\tilde{\mu}_{t},\tilde{a}_{t}){}+{}\lambda\log(h_{C}(\tilde{\mu}_{t{}+{}1})-L_{h}C_{n,t{}+{}1})$
@@
C.3.2. Mean-Field Deep Deterministic Policy Gradient (MF-DDPG)
In this section, we adopt the Deep Deterministic Policy Gradient (DDPG) algorithm Lillicrap et al. (2015) to the MFC. DDPG is a model-free actor-critic algorithm, hence, it can optimize Section 4 without the assumption of a differentiable simulator. However, it can not be applied directly to the problem because the Q-value for $\tilde{z}_{n,t}=(\thinspace_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})$ is ambiguous. The important insight here is that the value of a certain state $\thinspace{}_{n,t}$ of the environment reflects the whole population, i.e., the expected reward over the remainder of an episode for a given $\tilde{\mu}_{n,t}$ if every agent in every state $\thinspace{}_{n,t}$ chooses actions following $\pi_{n,t}$. In essence, the Q-value is a function of $\tilde{\mu}_{n,t}$ and $\pi_{n,t}$ and not $\tilde{z}_{n,t}$.
To overcome this issue, we introduce the lifted mean-field Markov decision process (MF-MDP) similarly to Carmona et al. (2019b); Gu et al. (2020, 2021a); Motte and Pham (2019); Pásztor et al. (2023). First, we rewrite the reward function as a function of the mean-field distribution and the policy, i.e.,
$\tilde{r}(\tilde{\mu}_{n,t},\pi_{n,t}){}={}\int_{\mathcal{S}}r(s,\tilde{\mu}_{n,t},\pi_{n,t}(s,\tilde{\mu}_{n,t}))\tilde{\mu}_{n,t}(ds).$
To simplify the notation, we overload the notation $\bm{\pi}_{n}^{\psi}{}={}(\bm{\pi}_{n}^{\psi},\eta^{\psi})$ as described in Section C.3.1. Then, we restate Section 4 as